The polynomial expressions 12y^3-5y^2-3y and 5(3y+1)(2y-3) share a common binomial factor?
What binomial factor do they share?

Found 2 solutions by ankor@dixie-net.com, dkppathak:
Answer by ankor@dixie-net.com(22740)

12y^3-5y^2-3y and 5(3y+1)(2y-3) share a common binomial factor?
What binomial factor do they share?
Factor 12y^3 - 5y^2 - 3y to:
y(3y + 1)(4y - 3)
:
(3y+1), is the common factor

Answer by dkppathak(439)

The polynomial expressions 12y^3-5y^2-3y and 5(3y+1)(2y-3) share a common binomial factor?
The polynomial expressions 12y^3-5y^2-3y =
y(12y^2-5y-3)=y(12y^2-9y+4y-3)
y(4y-3)(3y+1) and other is 5(3y+1)(2y-3)
common bionomial factors are (3y+1)
(3y+1) is common factor share




0